OpenStudy (loser66):

\[f(x,y)= \begin{cases}2~~~~~~if ~~0\leq x\leq 1; 0\leq y\leq 1; 0\leq x +y\leq 1\\0~~~~~~elsewhere\end{cases}\] find \(P (x\leq 3/4, y\leq 3/4)\) Please help me find the limits ALGEBRAICALLY, I know how to find it out by geometric way. However, it works only for uniform function like this. I would like to know how to find it out by algebraic way so that I can work on any function.

OpenStudy (loser66):

After using geometry, I got \[\int_0^{1/4}\int_0^{3/4}2dydx +\int_{1/4}^{3/4}\int_0^{1-x}2 dydx\] How to work on it without the pic?

OpenStudy (math&ing001):

What do you mean by "without the pic" ?

OpenStudy (loser66):



OpenStudy (loser66):

Hence, when drawing the pic, we can find the upper and lower limits to put on the integral. However, it works if the function is uniform only.

OpenStudy (loser66):

if the function is conditional one, we HAVE TO find the limits algebraically; pic doesn't work.

OpenStudy (math&ing001):

I think solving the equation x+y<=1 should give us limits for each case

OpenStudy (loser66):

Show me, please.

OpenStudy (math&ing001):

Like for the first case, take y<=3/4 so 1-x<= 3/4 => x<=1/4 That gives limits for first term

OpenStudy (math&ing001):

Wait I got that all wrong

OpenStudy (loser66):

Now, suppose we don't have the pic and we don't know that the function can be break into 2 terms like that. How to do?

OpenStudy (math&ing001):

Normally beginning with something like this should give you the same expression you got. \[\int\limits_{0}^{3/4}\int\limits_{0}^{3/4} 2 dx dy [x+y \le 1] = \int\limits_{0}^{3/4}\int\limits_{0}^{\min(3/4;x)}2dx dy\] I tried playing with it a little bit, but didn't get much...

OpenStudy (loser66):

:(

ganeshie8 (ganeshie8):

Are you saying that you want to figure out the bounds of a double integral w/o using a sketch of the region ?

OpenStudy (loser66):

Yes, Sir

ganeshie8 (ganeshie8):

that would be like searching for an object in the darkness w/o turning on the flashlight you're holding in your hand

ganeshie8 (ganeshie8):

Could you give me one valid reason for not using the sketch of the region of integration ?

ganeshie8 (ganeshie8):

`if the function is conditional one, we HAVE TO find the limits algebraically; pic doesn't work. ` why ?

OpenStudy (loser66):

My prof said that using picture as above works only for uniform function, for other functions, we have to find it out algebraically.

OpenStudy (loser66):

Let me post his way, it is so complicated so that I don't get it.

ganeshie8 (ganeshie8):

Okay..

OpenStudy (loser66):

Thanks for the link, I got it. :)

OpenStudy (loser66):

However, that is the method we use in cal3 with a fixed function, right? like our uniform function here.

ganeshie8 (ganeshie8):

I am still going through replies in that link...

OpenStudy (loser66):

That is 2 methods of integration. dy dx and dx dy. Both them rely on the pic.